OpenStudy (anonymous):

Solve for θ: cos θ + sqrt(3)*sin θ = 0

OpenStudy (anonymous):

\[\cos \theta + \sqrt{3}*\sin \theta = 0\]

OpenStudy (anonymous):

Never mind I got it:\[\cos \theta = -\sqrt{3}*\sin \theta\] \[\frac{ \cos \theta }{ \sin \theta }=-\sqrt{3}\] \[\theta=arccot(-\sqrt{3})\] \[\theta = \frac{ 5\pi }{ 6 }\]